### v1

# Data-Driven Goal Recognition Design for General Behavioral Agents

Abstract
Goal recognition design aims to make limited modifications to decision-making environments with the goal of making it easier to infer the goals of agents acting within those environments. Although various research efforts have been made in goal recognition design, existing approaches are computationally demanding and often assume that agents are (near-)optimal in their decision-making. To address these limitations, we introduce a data-driven approach to goal recognition design that can account for agents with general behavioral models. Following existing literature, we use worst-case distinctiveness (wcd) as a measure of the difficulty in inferring the goal of an agent in a decision-making environment. Our approach begins by training a machine learning model to predict the wcd for a given environment and the agent behavior model. We then propose a gradient-based optimization framework that accommodates various constraints to optimize decision-making environments for enhanced goal recognition. Through extensive simulations, we demonstrate that our approach outperforms existing methods in reducing wcd and enhancing runtime efficiency in conventional setups, and it also adapts to scenarios not previously covered in the literature, such as those involving flexible budget constraints, more complex environments, and suboptimal agent behavior. Moreover, we have conducted human-subject experiments which confirm that our method can create environments that facilitate efficient goal recognition from real-world human decision-makers.
1 Introduction
With the rapid advancement of artificial intelligence (AI), there has been a surge in interest in human-AI collaboration, aiming to synergize human and AI capabilities across various domains such as gaming, e-commerce, healthcare, industrial engineering, and workflow productivity. Designing AI agents to work alongside humans requires these agents to understand and infer human goals and intentions. While there has been abundant research in goal recognition (Ramírez and Geffner, 2010; Sukthankar et al., 2014) that aims to infer human goals by observing their actions, this work focuses on the goal recognition design problem (Keren et al., 2014), where one needs to identify how to modify decision-making environments to enable better goal recognition.
Our work is built on the goal recognition design problem formulated by Keren et al. (2014). In their work, they proposed the worst-case distinctiveness (wcd) metric, defined as the maximum number of decisions an agent can make without revealing its goal, to measure the difficulty of inferring the agent’s goal. They then aimed to modify the decision-making environment, through removing allowable actions from states, to optimize this measure. Since the introduction of this work, there have been several follow-up works that extend the formulation to deal with different settings, such as stochastic settings (Wayllace et al., 2016, 2017) and partial observability (Keren et al., 2016). More discussion can be found in the survey by Keren et al. (2021).
While there has been significant progress in goal recognition design, there are two main limitations in the literature. First, current approaches require evaluating the difficulty of goal recognition, i.e., worst-case distinctiveness (wcd), for a large number of potential modifications to the decision-making environment. This process is computationally demanding and limits the scalability of these approaches, especially in virtual domains where the number of possible environment configurations is large. Second, decision-making agents in goal recognition design are often assumed to be optimal. While there are efforts addressing suboptimal agents (Keren et al., 2015; Wayllace and Yeoh, 2022), these primarily focus on settings where there is a limited number of deviations from optimal decision-making. This assumption of optimality is limiting in scenarios involving human agents, who are known to often systematically deviate from optimal decision-making due to cognitive and informational constraints (Kahneman, 2003; O’Donoghue and Rabin, 1999).
To address these limitations, we propose a framework for data-driven goal recognition design with general agent behavior. We explicitly incorporate models of agent behavior into the optimization framework to better represent behavioral agents. To tackle the computational challenges, our approach leverages data-driven methods for goal recognition design. The core idea involves building a machine learning oracle that predicts the difficulty of goal recognition (e.g., wcd) given a decision-making environment and an agent’s behavioral model. This oracle is trained on datasets generated from simulations, allowing for the evaluation of wcd in any given environment and agent model. Such an approach significantly accelerates the evaluation of wcd during run-time optimization. Once the machine learning oracle is established, we apply gradient-based optimization to minimize wcd, employing Lagrangian relaxation to manage constraints on environment modifications.
To evaluate our data-driven goal recognition design framework, we conducted a series of simulations. Initially, these simulations were carried out in the standard setup in the grid world envrionment assuming optimal agent behavior. We show that our approach outperforms existing baselines in reducing worst-case distinctiveness (wcd) and demonstrates considerably better run-time efficiency. We then conducted additional simulations to showcase that our approach can generalize to settings beyond the capabilities of existing approaches in the literature. These include scenarios involving flexible budget constraints, more complex environments, and suboptimal agent behavior. Lastly, we have conducted human-subject studies demonstrating that our method can be leveraged to design environments that facilitate efficient goal recognition from real-world human decision-makers. The results highlight the potential of our approach to enable more efficient human-AI collaboration.
Contributions.
The main contributions of this work can be summarized as follows.
  • $\bullet$
    We propose a data-driven optimization framework for goal recognition design. This framework comprises a predictive model that estimates the wcd for a given environment and a model of agent behavior. Leveraging the differentiable property of the predictive model, we utilize a gradient-based optimization method for goal recognition design. The framework is flexible to accommodate various environments, design spaces, and agent models.
  • $\bullet$
    Through extensive simulations, we show that our framework not only outperforms existing approaches in goal recognition design within standard settings but also adapts to scenarios that existing approaches cannot handle. These scenarios include those with general optimization criteria, complex environments, and suboptimal agent behavior.
  • $\bullet$
    Through real-world human-subject experiments, we demonstrate that our approach can adapt to agent models trained on human behavior. Furthermore, our framework can create environments that enable more effective goal recognition with real-world human decision-makers.
2 Related Work
Our work contributes to the expanding field of human-AI collaboration. Recent research has indicated that optimizing AI alone is insufficient for maximizing the performance of human-AI teams (Bansal et al., 2021; Carroll et al., 2019). To develop truly collaborative AI agents, it is crucial to equip them with the ability to comprehend and predict the intentions and goals of their human counterparts. This challenge lies at the heart of goal recognition research (Kautz and others, 1991; Ramírez and Geffner, 2010; Hong, 2001; Sukthankar et al., 2014; Pereira et al., 2017; Masters et al., 2021). Our work specifically focuses on goal recognition design, an extension of goal recognition that includes modifying environments to better facilitate the process of recognizing goals.
Goal recognition design was originally formulated by Keren et al. (2014). Since this seminal work, numerous research efforts have expanded the concept to accommodate stochastic environments (Wayllace et al., 2017; Wayllace and Yeoh, 2022), different levels of observability (Keren et al., 2016; Wayllace et al., 2020), and a variety of design spaces (Mirsky et al., 2019). The studies most closely aligned with our approach focus on suboptimal agents (Keren et al., 2015; Wayllace and Yeoh, 2022). However, these studies characterize suboptimality by limiting deviations from the optimal policy, a method that may not adequately represent behavior of human agents, who frequently deviate systematically from optimal decision-making. Additionally, most existing work requires evaluating the difficulty of goal recognition in run time for numerous environmental modifications, which constrains the scalability of these methods. Our work sets itself apart by broadening the goal recognition design question to incorporate general agent behavioral models and by implementing a data-driven optimization approach.
From a technical standpoint, we adopt a data-driven optimization approach for goal recognition design. The use of data-driven tools in optimization has gained increasing prominence in the field of mechanism design, as evidenced by a range of recent studies (Dütting et al., 2019; Feng et al., 2018; Golowich et al., 2018; Curry et al., 2020; Kuo et al., 2020; Rahme et al., 2020; Peri et al., 2021; Cornelisse et al., 2022; Yu et al., 2023). In this regard, our work closely aligns with the studies by Yu and Ho (2022) and Yu et al. (2023), who explicitly encoded human behavior and responses within their mechanism designs for optimizing the information policy in environment and information design. While our work shares a similar motivation, our approach and problem formulation differ from theirs.
3 Problem Formulation and Methods
3.1 Problem Setting
We formulate the decision-making environment and behavioral agent models. We also explain the commonly-used metrics for evaluating goal recognition difficulty.
Decision-making environment.
We define the decision-making environment as a Markov Decision Process (MDP), represented by $W=\langle S,A,P,R\rangle$. Here, $S$ denotes the set of states, $A$ represents the set of agent actions, $P(s^{\prime}|s,a)$ is the transition probability from state $s$ to state $s^{\prime}$ upon taking action $a\in A$, and $R(s,a,s^{\prime})$ represents the bounded reward received after taking action $a$ in state $s$ and reaching state $s^{\prime}$. To emphasize the goal recognition aspects of the problem, we introduce a set of goal states $G\subseteq S$. These goal states are terminal; that is, $P(g|g,a)=1$ $\forall g\in G,a\in A$.
Models of behavioral agents.
We represent the agent’s decision-making policy in a general form $\Pi:S\times T\rightarrow A$. Specifically, for an agent with a decision-making policy $\pi\in\Pi$, the agent will execute the action $\pi(s,t)$ in state $s$ at time $t$. The agent is conceptualized as a planner $H:W\rightarrow\Pi$, where the input is an environment $w\in W$, and the output is a policy $\pi\in\Pi$. To illustrate our formulation, consider a agent parameterized by a time-variant discounting function $d(t)$. The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$. We can also represent an agent with present bias (O’Donoghue and Rabin, 1999) by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$. We would like to highlight that our approach not only accounts for standard analytical closed-form expression of agent behavior. It can also accounts for scenarios where an agent policy $\pi$ is a machine learning model, i.e., a neural network trained on human behavioral data.
Worst-case distinctiveness (wcd).
To evaluate the difficulty of goal recognition, we adhere to standard literature and focus on the measure of worst-case distinctiveness (wcd(Keren et al., 2014), defined as the maximum number of steps an agent can take before revealing its goal. To compute the wcd for a given suboptimal agent $h\in H$ in an environment $w\in W$, we evaluate the path for the agent to each goal and compute the number of actions from the initial state that are identical for every goal.
3.2 Goal Recognition Design with General Behavioral Agents
We formalize the goal recognition design problem with general behavioral agents. Given an environment $w$ and an agent with a behavior model $h$, we denote the worst-case distinctiveness of environment $w$ for agent $h$ as $\textit{wcd}(w,h)$. Each type of modification $1\leq i\leq N$ will incur a cost $c_{i}(w,w^{\prime})$ that must fall in budget constraint $B_{i}$. The objective of the goal recognition design problem is to alter the environment from $w$ to $w^{\prime}$ in a way that minimizes $\textit{wcd}(w^{\prime},h)$, while satisfying the constraint that the cost of the modifications does not exceed the budget.
$\underset{w^{\prime}}{\text{minimize}}$ $\textit{wcd}(w^{\prime},h)$ (1)
subject to $c_{i}(w,w^{\prime})\leq B_{i},\forall 1\leq i\leq N$
3.3 Our Proposed Method
Existing approaches in goal recognition design require evaluating wcd for a large number of potential modifications to the environments to identify the optimal modifications. Since evaluating wcd requires evaluating the agent policy for multiple goals, it presents significant computational overhead and limits the scalability of these approaches. To overcome this challenge, we propose leveraging machine learning to expedite run-time computations. The main idea, as summarized in Figure 1, is to first train a machine learning model that predicts wcd for any given pair of decision-making environment $w$ and agent behavioral model $h$. After obtaining this machine learning model and utilizing its differentiable property, we develop an optimization framework that applies gradient-based optimization methods to the Lagrangian relaxation of the constrained optimization formulation in (1).
(a) Predictive model for wcd.
(b) Optimization framework.
Figure 1: We first train a predictive model to predict wcd from simulated data. We then perform gradient-based optimization that leverages the predictive model to identify environment modifications that minimize wcd with a given model of agent behavior.
Predictive model for wcd.
To build the predictive model for wcd, we first curate a training dataset through simulations. For a given environment $w$ and agent behavioral model $h$, we can obtain $\textit{wcd}(w,h)$ by solving for the agent’s actions towards different goals. After collecting a training dataset, we train the predictive model using a convolutional neural network. Implementation details are included in the appendix.
Optimization procedure.
After obtaining the predictive model, we develop a gradient-based optimization framework. Note that this optimization framework strictly generalizes the existing literature in goal recognition design where the space of modifications are limited (e.g., usually limits to blocking an action in MDP) (Keren et al., 2021).
The first step involves transforming the constrained optimization problem in (1) into an unconstrained optimization problem using Lagrangian relaxation:
$\mathcal{L}=\textit{wcd}(w^{\prime},h)+\sum_{i}\lambda_{i}(c_{i}(w,w^{\prime})-B_{i}).$
We then perform gradient descent on the relaxed Lagrangian. As environment modifications are often discrete (e.g., whether to block a cell in the grid world), we apply a discrete gradient descent procedure. Specifically, at each step of gradient descent, we obtain a gradient, which is a vector indicating the suggested change magnitude for each element (such as a cell in the grid world). We then select the element with the highest gradient value and make the corresponding change. Note that some suggested modifications may not be valid; for instance, we cannot block a cell that is already blocked in the grid world. In such cases, we proceed to the element with the next highest gradient value, continuing this process until a valid modification is made. This modification procedure is repeated until the gradient descent converges.
Note that with this Lagrangian relaxation, we cannot directly set the budget, however, based on duality, a larger Lagrangian multiplier $\lambda$ corresponds to a smaller budget $B$, in the original constrained optimization formulation. In our experiments, instead of selecting different budgets directly, we choose varying Lagrangian multipliers and record the realized costs of the modifications.
4 Experiment Setup
4.1 Benchmark Domains
We utilize two benchmark domains. The first is the standard basic grid world environment, commonly used in goal recognition design literature. The second is the Overcooked-AI environment (Carroll et al., 2019), an complex environment with a richer set of environment modification. This environment is particularly relevant to the downstream implications of our work, namely in supporting human-AI collaboration.
(a) Grid world.
(b) Overcooked-AI.
Figure 2: The example showcases two benchmark environments. The first, displayed on the left, is a grid world environment. The agent starts at a position marked ’S’ and aims to reach one of the goal positions labeled ’G’. The agent must navigate through the grid, avoiding blocked cells marked with ’x’. The second environment, illustrated on the right, is an Overcooked-AI setting. In this scenario, the agent’s objective is to pick up ingredients and complete the cooking of their target recipe, which constitutes their goal.
4.1.1 Benchmark Domain: Grid world
In the grid world domain, agents navigate a grid with several potential goals. Take, for example, a grid world environment as illustrated in Figure 2(a). In this environment, an optimal agent starts at point ’S’ and aims for one of the goals. Spaces marked ’X’ are blocked, barring the agent’s passage. In this particular environment, the worst-case distinctiveness (wcd) is 0, as an optimal agent, following the shortest path to either goal, would reveal its intended goal on the first move due to non-overlapping paths. Our experiments are primarily focused on grid world environments with two goals for simplicity. However, our approach is applicable to scenarios with more than two goals.
Design space of environment modifications.
In the context of goal recognition design in grid world, the space environment modification is often limited to adding blocks to spaces (Keren et al., 2014), also called action removal in the literature. In our work, we broaden the design space for modifications to also consider the unblocking of existing blocked spaces as potential modifications.
4.1.2 Benchmark Domain: Overcooked-AI
We also conduct our evaluations on a more complex domain: Overcooked-AI [^1] https://github.com/HumanCompatibleAI/overcooked_ai. This environment is based on the popular multiplayer game Overcooked. In each game, players (agents) collaborate to prepare and deliver specific recipes. Since the goal of our approach is to enable efficient inference of agent goals, we focus on the special case with a single agent. The environment is represented as a grid (see Figure 2(b)) with each cell specifying the object that is placed at the cell. The objects may be a counter, a tomato, an onion, a pot, a dish, a serving point, or an empty space. The agent can only occupy empty spaces in the environment and cannot step on any other objects in the environment. The agent can carry the movable objects, i.e. onion, tomato, and dish, and drop them on any of the other non-space objects (counter, serving point, or pot). To navigate in the environment, the agent can move left, right, north, or south and maintain an orientation that is consistent with the last movement direction.
Goal recognition in Overcooked-AI.
The goal of the agent in Overcooked-AI is defined by the recipe it needs to complete. For instance, to prepare a soup with one onion and two tomatoes, the agent must pick up the ingredients, place them in the pot, and cook them. In the context of goal recognition, the objective is to deduce which recipe the agent is preparing. For the sake of simplicity in our experiments, we focus on scenarios with only two possible goals or recipes.
Design space of environment modifications.
A primary challenge in the Overcooked-AI environment is its considerably larger design space for environment modifications. The design space includes the addition, removal, or changing position of any object within the environment, subject to the modification being valid. This means an object cannot be added to a cell already occupied by another object, and an object that does not exist cannot be removed. The objective of conducting experiments in this domain is to assess whether our approach can be used to address more complex domains.
4.2 Baselines
We now describe the baseline approaches we use to compare against our approach.
  • $\bullet$
    Exhaustive search: This is the brute-force approach that evaluates wcd for all the environments on the search path until the minimum possible wcd is found. It is guaranteed to achieve minimum wcd. However, given the computation overhead, this approach is not applicable in most situations.
  • $\bullet$
    Pruned-Reduce (Keren et al., 2014): This baseline is specifically designed for settings where modifications are limited to action removal such as blocking a cell in grid world. It speeds up the exhaustive search and retains the optimal property. However, its scalability is still limited.
  • $\bullet$
    Greedy search using true wcd: This greedy search baseline finds the single environment modification that leads to the maximum reduction of wcd at each iteration. This approach requires to evaluate the wcd for all possible single environment modifications at each iteration.
  • $\bullet$
    Greedy search using predicted wcd: In addition to greedy search using true wcd, we leverage our predictive model for wcd and design another greedy baseline. This baseline finds the single environment modification that leads to the maximum reduction of predicted wcd at each iteration.
4.3 Models of Agent Behavior
In our experiments, to demonstrate that our approach works for different models of agent behavior, we have examined three types of agent behavior.
  • $\bullet$
    Optimal agent behavior. The first one is the standard optimal agent behavior. Conducting experiments with optimal agent behavior enables us to compare our approach with standard approaches in the literature, which are often developed under the optimal agent assumption.
  • $\bullet$
    Parameterized suboptimal agent behavior. We also consider an generalized behavior model that is captured by a time-variant discounting function $d(t)$, and the agent’s objective is to optimize a time discounted reward with the discounting factor for $t$ steps in the future being $d(t)$. The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$. We can also represent an agent with present bias (O’Donoghue and Rabin, 1999) by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
  • $\bullet$
    Data-driven agent behavior. We also address settings where the model of agent behavior is a machine learning model trained on human behavioral data.
5 Experiments
We have conducted extensive simulations to examine the performance of our approach. Moreover, we have conducted real-world human-subject experiments to assess whether our approach is effective with human decision-makers.
5.1 Simulations
In our simulations, we first examine how our approach compares with existing methods in standard setups found in the literature. We then showcase the generalizability of our approach by extending it to scenarios that go beyond the setups studied in the literature, including scenarios with more dynamic budget constraints, more complicated environments, and suboptimal agent behavior.
5.1.1 Standard setting
We first focus on settings within the grid world domain, where modifications are limited to blocking cells and agent behavior is assumed to be optimal. This scenario forms the standard setting for the majority of goal recognition design studies, as highlighted in Table 1 of the survey by Keren et al. (2021). The objective of this set of simulations is to enable comparisons with state-of-the-art methods.
In our experiments, the initial environment is generated randomly: we first randomly select the number of blocked cells from the range $[0,12]$, followed by the random allocation of the blocked cells’ positions. We also randomly determine the starting grid and two goal grids. Environments where the goals are not reachable from the starting grid are filtered out. Following this procedure, we randomly generate 200 environments and compare the average performance of our approach with that of baseline methods.
(a) $6\times 6$ grids
(b) $13\times 13$ grids
Figure 3: wcd reduction in a grid world when only blocking modifications are allowed. Exhaustive search and Pruned-Reduce are not included in the larger environments of (b) because they cannot complete the computation within an hour for a single environment.
We start our experiments with a grid world of size $6\times 6$. In this simplest scenario, our approach and all baselines exhibit similar performance in wcd reduction, as depicted in Figure 3(a). For runtime comparison, our method demonstrates a significant speed advantage over exhaustive search, performing on average 10 times faster. With the maximum allowed budget of 19 modifications, our approach required only 0.2 seconds, whereas the exhaustive search took approximately 2 seconds. [^2] Detailed runtime comparisons are included in the appendix. In this simplest setting, the runtime of our approach roughly matches that of the greedy method but is 10 times faster than exhaustive search. In all other settings, our method is several orders of magnitude faster than both exhaustive search and greedy using true wcd. Although the greedy baseline becomes much faster when utilizing our predicted model for wcd, our approach remains over 3 times faster than the greedy baseline with predicted wcd for large budgets. We then expanded the grid size to $13\times 13$. In this more complex setting, both exhaustive search and the Pruned-Reduce method were unable to complete computations within an hour for an instance, leading to their exclusion from the baselines. As illustrated in Figure 3(b), our approach surpasses the greedy baselines in wcd reduction and is 3 times faster than these baselines for large budgets.
5.1.2 Settings with flexible budget constraints
In the literature, most works focus on a single type of environment modification (e.g., blocking a cell). Given the flexibility of our optimization framework, we extend our approach to also include ’unblocking’ blocked cells as a possible environment update. In our simulations, we examine two common cases. In the shared budget setting, the total number of blocking and unblocking actions is bounded by a given shared budget. In the individual budget setting, the number of blocking and unblocking actions is bounded by different budgets. Specifically, we allow the number of blocking actions to be 5 times the number of unblocking actions (rounded to the nearest integer). Given that there are no established baselines for this setting in the literature, we compare our results against greedy baselines. The results, as shown in Figure 4, demonstrate the effectiveness of our approach in wcd reduction. Regarding runtime, our approach is several orders of magnitude faster than the greedy method with true wcd and at least 3 times faster than the greedy method with predicted wcd.
(a) Shared budget.
(b) Individual budget.
Figure 4: wcd reduction in settings with two types of modifications.
5.1.3 Complex domain and suboptimal agent behavior
We consider two additional extensions to the standard setting. In the first, we evaluate the performance of our approach in a more complex problem domain: Overcooked-AI. In the second, we return to the grid world but include scenarios with suboptimal agent behavior. Both extensions utilize a grid size of $6\times 6$. For the Overcooked-AI environment, we assume optimal agent behavior and aim to explore how our approach adapts to a much richer space of environment modifications. For the suboptimal human behavior, we employ the model described in Section 4.3, utilizing a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$ and set $k=8$. For both extensions, standard approaches such as exhaustive search and Prune-Reduce are either too slow or not applicable. Therefore, we compare our results with the greedy baselines. The results for both extensions, as illustrated in Figures 6 and 6, demonstrate that our approach adapts well to both settings. Regarding runtime, our approach is again several orders of magnitude faster than the greedy method with true wcd and at least 3 times faster than the greedy method using predicted wcd.
Figure 5: Overcooked-AI Environment. Assuming optimal agent behavior.
Figure 6: Grid world of size $6\times 6$. Incorporating suboptimal agent behavior.
5.2 Real-World Human-Subject Experiments
In our simulations, we demonstrate that our approach consistently outperforms baseline methods in terms of wcd reduction and offers greater efficiency in runtime. To assess the applicability of our approach in settings where humans are the decision-makers, we conducted two sets of experiments involving human subjects. In the first experiment, our goal is to collect human behavioral data within our decision-making environments. Utilizing this data, we employ imitation learning to develop a model that accurately represents human behavior. This model is then integrated as the agent behavior model within our approach. In our second experiment, we aim to evaluate whether our approach indeed leads to environments that facilitate more effective goal recognition by human decision-makers. These experiments are approved by the Institutional Review Board (IRB) of our institution.
5.2.1 Experiment 1: Collection of human behavioral data
In our experiment, we recruited participants to play 15 navigation games within a $6\times 6$ grid world. In each game, participants were tasked with navigating from a start position to a designated goal. At each time step, they could choose to move in one of four directions: Up, Down, Right, or Left. A game concluded when the participant reached the goal. The environments for these games were generated similarly to our simulations, with start positions, goal positions, and block positions all being randomly determined. Our objective with this setup was to leverage the collected data to develop a data-driven model of human behavior. We recruited 200 workers from Amazon Mechanical Turk for this study, paying an average hourly rate of approximately \$16 per worker.
Figure 7: The interface of our human subject experiment.
Table 1: The prediction accuracy for human behavior assuming optimal behavior and using data-driven model in Experiment 1.
Training Accuracy Validation Accuracy Testing Accuracy
Assuming Optimal Behavior 0.7266 0.6964 0.7131
Data-Driven Model 0.9189 0.8136 0.8422
Learning models of human behavior.
The collected human data were divided into training, validation, and testing sets: the training set included data from 160 workers, comprising approximately 70,000 instances of user decisions, while the validation and testing sets each contained data from 20 workers, amounting to around 8,800 instances of user decisions each. We employed a 4-layer Multilayer Perceptron (MLP) model for training the model. The input to the model is the current environment layout, and its output predicts the next human action. We fine-tuned the hyperparameters, including learning rate, hidden layer size, and L2 penalty, based on the performance on the validation dataset.
We compared the performance of our learned model against a model that assumes optimal agent behavior, which is defined as taking the shortest path to the goal. The training, validation, and test accuracies of both models are presented in Table 1. These results clearly indicate that human behavior deviates significantly from optimality. This deviation underscores the importance of incorporating a realistic model of human behavior in goal recognition design, particularly when humans are the decision-makers.
5.2.2 Experiment 2: Evaluating goal recognition design
We next evaluate the performance of our approach that incorporates the data-driven model of human behavior from experiment 1. To do this, we randomly generate 30 initial environments using the same setup as in our simulations. These environments are then updated according to four different methods, all operating within a modification budget of 20.
  • $\bullet$
    Initial: No updates to the environment.
  • $\bullet$
    Greedy: Greedy baseline using predicted wcd from the data-driven human behavior model.
  • $\bullet$
    Proposed (opt-bhvr): Our proposed approach when assuming the agent is following optimal behavior (i.e., picking one of the shortest paths towards the goal).
  • $\bullet$
    Proposed (data-driven): Our proposed approach using predicted wcd from the data-driven model.
We recruited 200 workers from Amazon Mechanical Turk. Each worker was randomly assigned to one of the four treatments specified above, with the distinction between treatments being the environments presented to them. Workers were randomly assigned a goal for each environment and tasked with navigating to reach it. After collecting their series of actions aimed at reaching the goal, we analyzed this behavioral data to assess the effectiveness of goal recognition design for each approach.
Comparing average overlapping path.
To evaluate the effectiveness of each goal recognition design approach, we first measure the average overlapping path towards each of the two goals. Specifically, each recruited worker is exposed to all 30 environments in their assigned treatment and is instructed to reach one of two randomly selected goals. For each treatment, we calculate the number of overlapping actions for every pair of workers assigned to different goals within each environment. We then compute the average of these numbers, yielding the average overlapping path as our performance measure. This metric reflects the difficulty of inferring agent goals for the envioronment and serves as an empirical proxy for wcd. With this metric, we vary the initial $k$ portion of actions taken by the workers in our analysis and observe how the average overlapping path changes with the amount of observed actions, denoted by $k$. As depicted in Figure 10(c), our approach, when integrated with data-driven models, leads to environments that facilitate the easier identification of the humans’ intended goal.
Comparing the accuracy of goal inference.
We next measure the accuracy of inferring human goals based on partial observations, utilizing an off-the-shelf Bayesian inference algorithm [^3] We assume an initial uniform prior distribution. As the agent takes actions, we update the posterior belief over goals based on the likelihood probability of the data-driven behavior model prediction. . Specifically, for each worker, we reveal the first $k$ portion of the worker’s actions to the inference algorithm to infer the worker’s goal. This enables us to compute the average inference accuracy. The results, as shown in Figure 8(b), demonstrate that our approach leads to environments that are easier for goal recognition.
(a) Averaging overlapping path (the lower the better.)
(b) Bayesian goal inference (the higher the better).
Figure 8: Comparing different goal recognition design approaches in our human subject experiments. Our approach coupled with data-driven models is shown to generate environments that enable the most effective goal recognition with real human decision makers.
6 Conclusion
Effective collaboration between humans and AI hinges on the AI’s ability to infer intentions and goals of humans. In this study, we work on the problem of goal recognition design, aiming to update the decision-making environments to make it easier for the AI agents to perform goal recognition. By addressing the key limitations in the existing literature, notably the computational demand and the assumption of optimal agent behavior, we have developed a data-driven optimization framework that is both efficient and scalable. Moreover, our approach is capable of integrating general models of various agent behavior, moving beyond the traditional assumption of optimality and making it more applicable to real-world scenarios involving human agents.
Future work of our study includes extending our framework to more complex environments, particularly those that are dynamic and partially observable. Given our framework is adaptable to different forms of agent behavior, understanding how different patterns of agent suboptimality impacts the difficulty of goal recognition could lead to useful insights for enabling better human-AI collaboration.
Appendix A Experiment Details and Additional Results
In this section, we provide details of the experiments that have not been included in the main paper due to space constraints. We also include and discuss additional experiment results.
A.1 Implementation Details
In this subsection, we discuss the details of our implementation for training the predictive model for wcd. To train a CNN-based prediction model, we generated an extensive training dataset by solving wcd for a large number of randomly selected environments, given a specific agent model. For creating a predictive model for environments with a grid size of $N$ by $N$, we utilized a custom Convolutional Neural Network (CNN) architecture. This architecture is tailored to process input data, $w_{i}$, consisting of $k$ $N\times N$ channels, where $k$ represents the number of potential objects in the environment.
The architecture comprises two distinct blocks: the initial block contains three successive convolutional layers, each followed by batch normalization and a rectified linear unit (ReLU) activation function. The subsequent block introduces pooling operations and additional convolutional layers for more sophisticated feature extraction. After the convolutional blocks, the architecture includes fully connected layers, leading to a single output unit. The design integrates Leaky ReLU, ReLU, and dropout layers to add non-linearity and provide regularization. This flexible architecture can adapt to tasks with multi-channel input data, allowing for modifications based on the specific needs of the task and dataset characteristics. Batch normalization and dropout are incorporated to promote training stability and prevent overfitting. We employed the Adam optimizer and Mean Squared Error (MSE) loss for training the models, specifically focusing on settings involving optimal human agents.
To create the dataset to train the model, we randomly generated $150$K grid designs and computed the corresponding wcd. This dataset was then split into the training dataset ($80\%$) and validation dataset ($20\%$). For our experiments reported in Section 5 with $13\times 13$ grid size and optimal agent behavior, we have achieved an MSE value around 0.18 for both training and validation errors. The results are shown in Figure 9. We have also examined a variety of setups, and generally we can reach small errors for the predictive model that can enable the optimization.
Figure 9: Training loss and validation errors for the predictive model for wcd with $13\times 13$ grid world and optimal human behavior.
A.2 Results for Run Time Comparisons
Due to space constraints and also that the results align with one would expect, we do not include the run time details for different approaches in the main text. We include the results here for completeness. Overall, as shown in Figure 10, approaches that leverage the predictive model for wcd are orders of magnitude faster than methods that require to evaluate wcd during run-time. Note that the y-axis of the figure is in log scale so the difference is in at least two orders of magnitude.
(a) Standard Setup (small)
(b) Standard Setup (large)
(c) Suboptimal
(d) Overcooked-AI
Figure 10: Run time taken by each method in the different experimental conditions. In the standard setup with a small grid size, all methods except exhaustive complete within less than half of a second but only our approach scales with a larger grid size and other more complex configurations.
Appendix B Additional Experiment Results
In this section, we report the additional experiment results that are not included in the main text due to space constraints. In Section 5.1.2 of the main paper, we provided details of our performance with a large grid size. In a smaller grid world, our approach achieves comparable performance to the baselines but it significantly outperforms them in larger grid sizes as shown in Figure 11.
(a) Individual budget
(b) Shared budget
Figure 11: wcd reduction in smaller ($6\times 6$) gridworld with optimal agent behavior.
Appendix C Simulation Details
To generate our training and evaluation datasets, we randomly generated environments and kept environments that are valid, e.g., the goals are reachable, and that the objects don’t overlap. Below we provide more details for environment generation for specified grid sizes for both overcooked and grid-world domains.
C.0.1 Gridworld
In a grid world, a valid environment includes a starting position, blocked cells, and two goal positions. The starting position is randomly placed in the first column, and the goal positions are randomly placed in the last two columns. The number of blocked positions in each grid is randomly selected from a range of 00 to 2×2$\times$2 × the grid width. We discard any assignments that make the goals unreachable. For experiments with suboptimal behavior, we also randomly assigned small subgoal rewards to unblocked cells that the agent would collect on its way to the goal state. The two goal states assigned a large reward that is 10 times the largest subgoal reward.
C.0.2 Overcooked-AI
In Overcooked-AI, a valid environment includes one pot, one tomato source, one onion source, one dish source, one serving point, no open spaces at the border, and any number of open spaces and blocked cells. All objects must be reachable from the agent’s randomly assigned starting position, with the number and positions of blocked cells assigned randomly. The agent is randomly assigned any two possible goals: three tomato soups, three onion soups, or a mixed soup. Each goal has the same randomly assigned reward value. Suboptimal behavior is modeled by assigning small random subgoal rewards when adding ingredients to the pot.
Appendix D Details of Human-Subject Experiment
Lastly, we include more information about our human-subject experiments. In the human-subject experiment, each worker is asked to play a navigation game in $6\times 6$ grid world environments. The task interface is shown in Figure 12.
Figure 12: Human-Subject experiment interface.
We have recruited 200 workers from Amazon Mechanical Turk in total, and Table 2 contains the demographic information of the workers.
Table 2: Demographic information of the participants in our experiment.
Group Category Number
Age 20 to 29 84
30 to 39 76
40 to 49 26
50 or older 14
Gender Female 89
Male 110
Other 1
Race / Ethnicity Caucasian 175
Black or African-American 8
American Indian/Alaskan Native 3
Asian or Asian-American 8
Spanish/Hispanic 1
Other 5
Education High school degree 5
Some college credit, no degree 5
Associate’s degree 4
Bachelor’s degree 135
Graduate’s degree 49
Other 2